what is the mass of sucrose in its 1L solution which is isotonic with 6.6 x 10^-3 kg/l of urea 
 

Grade:8

1 Answers

Shweta Ahire
5 years ago
6.6 × 10^-3 ×10^3 gm of urea present in 1lit 
So mole of urea =6.6 / 60 = 0.11
In Urea (NH2CONH2) ,Neutrons = 0.11× 6.022 ×10^23 × 28(total neutrons ) = 18.54 × 10^23
 
Isotonic  means same no. Of neutrons so sucrose also have 18.54 ×10^23 neutrons 
So mole of sucrose (C12H22O11) = (18.54 × 10^23)/(6 ×10^23 × 160).......Where  (160 = 12×6+8×11)
So moles of sucrose = 0.019
So mass of sucrose = mole × molecular weight 
                                    = 0.019 × 342
                                    = 6.604 gm